OpenStudy (anonymous):

can someone help me find what i did wrong and possibly walk me thru the rest so i have a better understanding? -(1+7x)-6(-7-x)=36 -1 +7x+42+6x=36.......7x+6x=13x & -1+42=41 13x + 41 = 36 -36 -36 thats where i get stuck and know i did something wrong but not sure what

OpenStudy (texaschic101):

-(1 + 7x) -6(-7-x) = 36 -1 - 7x + 42 + 6x = 36 -x + 41 = 36 -x = 36 - 41 -x = - 5 (multiply by -1 to make x positive) x = 5 check... -(1 + 7x) - 6(-7-x) = 36 -(1 + 7(5) - 6(-7-5) = 36 -(1 + 35) - 6(-12) = 36 -36 + 72 = 36 36 = 36 (correct) any questions ?

OpenStudy (anonymous):

not sure how you got the 41

OpenStudy (texaschic101):

I combined like terms -1 + 42 = 41

OpenStudy (texaschic101):

do you see it ?

OpenStudy (anonymous):

duh! these things are always getting me mixed up in one way or another. thank you

OpenStudy (texaschic101):

it is easy to get confused on these problems...that is why I always check them

OpenStudy (texaschic101):

you are very welcome...glad i could help